\begin{thebibliography}{10}
\csname url@rmstyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand\BIBentrySTDinterwordspacing{\spaceskip=0pt\relax}
\providecommand\BIBentryALTinterwordstretchfactor{4}
\providecommand\BIBentryALTinterwordspacing{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand\BIBforeignlanguage[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\else
\language=\csname l@#1\endcsname
\fi
#2}}

\bibitem{Avg:ub}
S.~V. Avgustinovich, ``On a property of perfect binary codes,''
  in \emph{Operations Research and Discrete Analysis},
  ser. \emph{Math. Appl.},
  A.~D. Korshunov, Ed.\hskip 1em plus 0.5em minus 0.4em\relax Kluwer,
  1997, vol. 391, pp. 13--15,
  translated from \em
  \href{http://www.math.nsc.ru/publishing/DAOR/daor.html}{Diskretn. Anal.
  Issled. Oper.}\em , 2(1) (1995), 4--6.

\bibitem{AvgHedSol:class}
S.~V. Avgustinovich, O.~Heden, and F.~I. Solov'eva, ``The classification of
  some perfect codes,'' \emph{\href{http://www.springerlink.com/content/100256/}{Des. Codes Cryptography}}, vol.~31, no.~3, pp.
  313--318, 2004, DOI:
  \href{http://dx.doi.org/10.1023/B:DESI.0000015891.01562.c1}{10.1023/B:DESI.0%
000015891.01562.c1}.

\bibitem{AvgSol:alpha-comp}
S.~V. Avgustinovich and F.~I. Solov'eva, ``Construction of perfect binary codes
  by sequential shifts of $\tilde\alpha$-components,''
  \emph{\href{http://www.springerlink.com/content/0032-9460}{Probl. Inf.
  Transm.}}, vol.~33, no.~3, pp. 202--207, 1997,
    translated from \emph{Probl. Peredachi Inf.} 33(3) (1997), 15--21.

\bibitem{EV:94}
T.~Etzion and A.~Vardy, ``Perfect binary codes: Constructions, properties and
  enumeration,''
  \emph{\href{http://ieeexplore.ieee.org/xpl/RecentIssue.jsp?punumber=18}{IEEE
  Trans. Inf. Theory}}, vol.~40, no.~3, pp. 754--763, 1994, DOI:
  \href{http://dx.doi.org/10.1109/18.335887}{10.1109/18.335887}.

\bibitem{KOP}
P.~Kaski, P.~R.~J. {\"O}sterg{\aa}rd, and O.~Pottonen, ``The {S}teiner
  quadruple systems of order $16$,''
  \emph{\href{http://www.sciencedirect.com/science/journal/00973165}{J. Comb.
  Theory, Ser.~A}}, vol. 113, no.~8, pp. 1764--1770, 2006, DOI:
  \href{http://dx.doi.org/10.1016/j.jcta.2006.03.017}{10.1016/j.jcta.2006.03.0%
17}.

\bibitem{Kro:quas_l_b}
D.~S. Krotov, ``Lower estimates for the number of $m$-quasigroups of order $4$
  and for the number of perfect binary codes,''
  \emph{\href{http://www.math.nsc.ru/publishing/DAOR/daor.html}{Diskretn. Anal.
  Issled. Oper.}, Ser. 1}, vol.~7, no.~2, pp. 47--53, 2000, in Russian.

\bibitem{KroPot:Lyap}
D.~S. Krotov and V.~N. Potapov, ``On the reconstruction of $n$-quasigroups of
  order $4$ and the upper bounds on their number,'' in \emph{Proc. the
  {C}onference Devoted to the 90th Anniversary of {A}lexei A. {L}yapunov},
  Novosibirsk, Russia, Oct. 2001, pp. 323--327, available at
  \url{http://www.sbras.ru/ws/Lyap2001/2363}.

\bibitem{MWS}
F.~J. MacWilliams and N.~J.~A. Sloane, \emph{The Theory of Error-Correcting
  Codes}.\hskip 1em plus 0.5em minus 0.4em\relax Amsterdam, Netherlands: North
  Holland, 1977.

\bibitem{Mal:lb}
S.~A. Malyugin, ``On a lower bound on the number of perfect binary codes,''
  \emph{\href{http://www.sciencedirect.com/science/journal/0166218X}{Discrete
  Appl. Math.}}, vol. 135, no. 1-3, pp. 157--160, 2004, translated from \em
  \href{http://www.math.nsc.ru/publishing/DAOR/daor.html}{Diskretn. Anal.
  Issled. Oper.}, Ser. 1\em , 6(4) (1999) 44--48. DOI:
  \href{http://dx.doi.org/10.1016/S0166-218X(02)00302-5}{10.1016/S0166-218X(02%
)00302-5}.

\newpage
\bibitem{Phelps84}
K.~T. Phelps, ``A general product construction for error correcting codes,''
  \emph{SIAM J. Algebraic Discrete Methods}, vol.~5, no.~2, pp. 224--228, 1984.

\bibitem{PheVil:RankKernel}
K.~T. Phelps and M.~Villanueva, ``On perfect codes: Rank and kernel,''
  \emph{\href{http://www.springerlink.com/content/100256/}{Des. Codes Cryptography}}, vol.~27, no.~3, pp. 183--194, 2002, DOI:
  \href{http://dx.doi.org/10.1023/A:1019936019517}{10.1023/A:1019936019517}.

\bibitem{PotKro:asymp}
V.~N. Potapov and D.~S. Krotov, ``Asymptotics for the number of $n$-quasigroups
  of order $4$,''
  \emph{\href{http://www.springerlink.com/content/0037-4466}{Sib. Math. J.}},
  vol.~47, no.~4, pp. 720--731, 2006, DOI:
  \href{http://dx.doi.org/10.1007/s11202-006-0083-9}{10.1007/s11202-006-0083-9},
  translated from \href{http://math.nsc.ru/smz/}{\emph{Sib. Mat. Zh.}} 47(4) (2006),
  873--887.
  ArXiv:~\href{http://arxiv.org/abs/math/0605104}{math/0605104}

\bibitem{Vas:nongroup_perfect}
Y.~L. Vasil'ev, ``On nongroup close-packed codes,'' in \emph{Problemy
  Kibernetiki}, 1962, vol.~8, pp. 337--339, in Russian, English translation in
  \emph{Probleme der Kybernetik}, 8 (1965), 92--95.

\bibitem{ZinLob:2000}
V.~A. Zinoviev and A.~Lobstein, ``On generalized concatenated constructions of
  perfect binary nonlinear codes,''
  \emph{\href{http://www.springerlink.com/content/0032-9460}{Probl. Inf.
  Transm.}}, vol.~36, no.~4, pp. 336--348, 2000,
  translated from \emph{Probl. Peredachi Inf.} 36(4) (2000), 59--73.

\bibitem{ZZ:2006:l16r14}
V.~A. Zinoviev and D.~V. Zinoviev, ``Binary extended perfect codes of length 16
  and rank 14,''
  \emph{\href{http://www.springerlink.com/content/0032-9460}{Probl. Inf.
  Transm.}}, vol.~42, no.~2, pp. 123--138, 2006, DOI:
  \href{http://dx.doi.org/10.1134/S0032946006020062}{10.1134/S0032946006020062},
  translated from \emph{Probl. Peredachi Inf.} 42(2) (2006), 63--80.

\textbf{\underline{New references:}}

\bibitem{KroPot:n-ary}
D.\,S.\,Krotov, V.\,N.\,Potapov,
\href{http://dx.doi.org/10.1137/070697331}{$n$-Ary quasigroups of order 4},
\href{http://epubs.siam.org/SIDMA}{\emph{SIAM J. Discrete Math.}}
\href{http://siamdl.aip.org/dbt/dbt.jsp?KEY=SJDMEC&Volume=23&Issue=2}%
{23(2)}
2009, 561--570,
{  \footnotesize
  DOI: \href{http://dx.doi.org/10.1137/070697331}{10.1137/070697331}
  ArXiv:\,\href{http://arxiv.org/abs/math/0701519}{math/0701519}
}

\bibitem{OstPot:16}
P.\,R.\,J.\,\"Osterg{\aa}rd,  O.\,Pottonen,
{The perfect binary one-error-correcting codes of length $15$: part I -- classification},
{\emph{IEEE Trans. Inform. Theory}}
\href{http://ieeexplore.ieee.org/xpl/tocresult.jsp?isnumber=5238716&isYear=2009}%
{55(10)} 2009, 4657--4660,
{  \footnotesize
  DOI: \href{http://dx.doi.org/10.1109/TIT.2009.2027525}{10.1109/TIT.2009.2027525}
  ArXiv:\,\href{http://arxiv.org/abs/0806.2513}{0806.2513}
}

\bibitem{OPP:16}
P.\,R.\,J.\,\"Osterg{\aa}rd,  O.\,Pottonen, K.\,T.\,Phelps,
{The perfect binary one-error-correcting codes of length $15$: part II -- properties},
\emph{submitted for publication},
  ArXiv:\,\href{http://arxiv.org/abs/0903.2749}{0903.2749}

\end{thebibliography}

\documentclass[letterpaper,journal,twocolumn,twoside,final,9pt]{IEEEtran}
\usepackage{amssymb,amsfonts}

\providecommand\href[2]{#2}
\newtheorem {theorem}{Theorem}
\newtheorem {corollary}{Corollary}

\newtheorem {proposition}{Proposition}
\newtheorem {remark}{Remark}
\newtheorem {constr}{Construction}
\newtheorem {conjecture}{Conjecture}

\newcommand\F[1][n]{F^{#1}_{\scriptscriptstyle\rm E\hspace{-0.1ex}v}}
\newcommand\FF[1][n]{F^{#1}_{\scriptscriptstyle\rm O\hspace{-0.1ex}d}}

\def\eqdf{\triangleq}

\begin{document}
\def\hrefDAOR#1{\href{http://math.nsc.ru/conference/DAOR'04/main.html}{#1}}
\title{On the Number of $1$-Perfect Binary Codes:\\ a Lower Bound%
\thanks{
This is author's version of the correspondence in the
\href{http://ieeexplore.ieee.org/xpl/RecentIssue.jsp?punumber=18}{IEEE Transactions on Information Theory}
54(4) 2008, 1760-1765,
Digital Object Identifier
\href{http://dx.doi.org/10.1109/TIT.2008.917692}{10.1109/TIT.2008.917692},
\copyright 2008 IEEE.
Personal use of the material of the correspondence is permitted.
However, permission to reprint/republish
the material for advertising or promotional
purposes or for creating new collective works
for resale or redistribution to servers or lists,
or to reuse any copyrighted component of this work
in other works must be obtained from the IEEE.}
\thanks{The material in this correspondence was presented in part at the
10th International Workshop on Algebraic and Combinatorial Coding Theory
ACCT-10, Zvenigorod, Russia, September 2006 
and, in a brief form,
at the Russian conference
``\hrefDAOR{{D}iscrete Analysis and} \hrefDAOR{Operations Research}''
\hrefDAOR{{DAOR}'2004},
Novosibirsk, Russia, June-July 2004, \href{http://www.math.nsc.ru/conference/DAOR'04/daor4.pdf}{p95}.}
}
\author{\href{http://arxiv.org/find/grp_math/1/au:+Krotov_Denis/0/1/0/all/0/1}{Denis~S.~Krotov} and
\href{http://arxiv.org/find/grp_math/1/au:+Avgustinovich_Sergey/0/1/0/all/0/1}{Sergey~V.~Avgustinovich}
\thanks{The authors are with the Sobolev Institute of Mathematics,
pr-t Ak. Koptyuga 4, Novosibirsk, 630090, Russia
(e-mail: krotov@math.nsc.ru, avgust@math.nsc.ru)}
 }
\maketitle
\begin{abstract}
We present a construction of $1$-perfect binary codes,
which gives a new lower bound on the number of such codes.
We conjecture that this lower bound is asymptotically tight.

\emph{Index terms}---{automorphism, lower bound, perfect binary codes}
\end{abstract}

\newcommand{\EExpLog}[2][0ex]{{{\raisebox{#1}{$2$}}^{\frac{n}{#2}-{\log}\frac{n}{#2}-1}}}
\newcommand{\EExp}[2][0ex]{{{\raisebox{#1}{$2$}}^{\frac{n}{#2}-1}}}
\newcommand{\EexpLog}[2][-0.2ex]{{{\raisebox{#1}{$\scriptstyle 2$}}^{\frac{n}{#2}-{\log}\frac{n}{#2}-1}}}
\newcommand{\Eexp}[2][-0.2ex]{{{\raisebox{#1}{$\scriptstyle 2$}}^{\frac{n}{#2}-1}}}
\newcommand{\ExpHam}[2][0ex]{2^{\EexpLog[-0.2ex]{#2}}}

\section{ Introduction}
The paper is devoted to the problem of enumeration of $1$-perfect binary codes.
Such codes, as any optimal codes,
are extremal objects of the theory of error-correcting codes.
In addition, perfect codes are a special type of combinatorial configurations.
The construction of $1$-perfect binary codes presented in the current paper gives
the most powerful known class of such codes and leads to a lower bound on their number.

The first known construction \cite{Vas:nongroup_perfect}
of nonlinear $1$-perfect binary codes gives the lower bound
\begin{eqnarray} \nonumber
B(n-1)&\geq&\ExpHam{2}\cdot\ExpHam{4}\cdot\ExpHam{8}\cdot      \dots
\end{eqnarray}
on the number $B(n-1)$ of $1$-perfect codes of length $n-1=2^m-1$ (here
and in what follows ${\log}$ means ${\log}_2$).
This bound was improved in \cite{AvgSol:alpha-comp} and \cite{Mal:lb},
where some useful ideas exploited 
in this paper were proposed.
The best known lower bound \cite{Kro:quas_l_b} is
\begin{eqnarray} \label{eq:0}
B(n-1)&\geq&\ExpHam{2}\cdot 3^{\Eexp{4}}\cdot \ExpHam{4}.
\end{eqnarray}
The result of \cite{Kro:quas_l_b} was formulated in the terms of a partial case \cite{Phelps84}
of the generalized concatenation construction (see, e.g., \cite{ZinLob:2000}),
which allows to construct $1$-perfect binary codes from distance $2$
$q$-ary MDS codes (only the case $q=4$ is useful for the lower bound),
or $n$-ary quasigroups (of order $4$).
The lower bound on the number of
$n$-ary quasigroups of order $4$ given in \cite{Kro:quas_l_b}
is asymptotically tight \cite{KroPot:Lyap,PotKro:asymp};
therefore, such a way to evaluate the number
of $1$-perfect binary codes has been exhausted.

The best known upper bound \cite{Avg:ub} on the number of $1$-perfect binary codes
is
$
2^{2^{n-(3/2){\log} n+\log{\log} (e n)}}.
$

The local-automorphism method presented in this work is a further development
of the methods \cite{Vas:nongroup_perfect,AvgSol:alpha-comp,Mal:lb,Kro:quas_l_b}.
Since there is one-to-one correspondence between $1$-perfect binary codes
and extended $1$-perfect binary codes, the results are formulated in terms of
extended $1$-perfect binary codes.

\section{ Preliminaries}
Let $F^n$ ($\F$, $\FF$) be the set of the binary $n$-words
(with even or odd number of ones, respectively) with the Hamming distance $d$
and mod~$2$ coordinate-wise addition. Given $\bar x \in F^n$, put $wt(\bar x) \eqdf d(\bar x,\bar 0)$.
The {\it neighborhood} of $S\subseteq F^n$ is the set
$\Omega(S)\eqdf \bigcup_{\bar x\in S}\Omega(\bar x)$ where
$\Omega(\bar x)\eqdf \{\bar y\in F^n \,|\,  d(\bar y,\bar x)=1\}$.

A set $C\subseteq F^n$ is called a \emph{distance $d$ code} (\emph{of length $n$})
{{if}} the Hamming distance between any two different words in $C$ is not less than $d$.
An {\it extended $1$-perfect code\,} is a set $C\subseteq \F$
such that the neighborhoods of the words of $C$
are pairwise disjoint and $\Omega(C)=\FF$.
It follows that $C$ is a distance $4$ code
of cardinality $|C|=|\FF|/n=2^{n-{\log} n-1}$ and $n$ is a power of $2$.
On the other hand, if the code $C\subset \F$ has distance $4$ and $|C|=2^{n-{\log} n-1}$, then, obviously, $C$ is an extended $1$-perfect code.
In what follows we assume $n=2^m\geq 16$.

The following formulas define some useful sets $V^t,A^t\subset\F$ and
give a representation of the extended Hamming code:
\begin{eqnarray}
V^t & \eqdf & \{(\bar{v},\bar v,0,\ldots ,0)\in F^n\,|\,\bar v\in \F[2^{m-t}]\},
\\ \nonumber&&
 t=1,\ldots,m-1
\\
 A^1 & \eqdf & V^1
\nonumber \\
A^{t} & \eqdf & V^t + A^{t-1} = \bigcup_{\bar r\in V^{t}}(\bar r+A^{t-1}), \label{eq:At}
\\ \nonumber&&
 t=2,\ldots,m-1
 \\
H & \eqdf & A^{m-1}. \label{eq:1}
\end{eqnarray}
The following is straightforward:
\begin{proposition}
The set $H$ defined by {\rm (\ref{eq:1})} is a linear extended $1$-perfect code,
i.\,e., the extended Hamming code (see e.\,g. {\rm\cite[\S1.7]{MWS}}),
which is the only linear extended $1$-perfect code,
up to coordinate permutation.
\end{proposition}

We say that a set $G\subset \F$ is a \emph{component of order $t\in\{1,\ldots,m-1\}$}
{{if}} $|G|=|A^t|$ and $\Omega(G)=\Omega(A^{t})$.
Let $t\in\{1,\ldots,m-1\}$ and $\bar \mu\in F^{2^{m-t}}\times {0}^{n-2^{m-t}}$;
we say that a set $M\subset \F$
is a \emph{$\bar \mu$-component} (of order $t$)
{{if}} $M=G+\bar \mu$ for some component $G$ of order $t$.

Let ${\rm Aut}(F^n)$ denote the group of isometries $F^n$
(it coincides with the automorphism group of the distance-one graph of $F^n$).
It is known that each isometry $g\in {\rm Aut}(F^n)$ has a unique
representation $g(\cdot) = \bar v+\pi(\cdot)$ where $\bar v$ is a \emph{shift vector}
 from $F^n$ and $\pi$ is a coordinate permutation.
If $\pi=Id$, i.\,e., $g(\cdot) = \bar v+\cdot$, then the isometry $g$ is called a \emph{translation}.
If $S\subseteq F^n$, then ${\rm Aut}({S})$ is the group of isometries $g$ of $F^n$ such that $g(S)=S$.
For a collection ${\mathbf{S}}=\{S_1,\ldots ,S_l\}$ of subsets of $F^n$,
by ${\rm Aut}(\mathbf{S})$ denote the group of isometries $g$ of $F^n$ such that
for each $S\in\mathbf{S}$ the set $g(S)$ is also in $\mathbf{S}$.
In what follows we use calligraphic letters to denote subgroups of ${\rm Aut}(F^n)$.
Put $${\cal A}^t \eqdf {\rm Aut}(\Omega(A^{t}))$$
where $A^t$ is specified by (\ref{eq:At}).

\begin{proposition}\label{pro:1} {\rm (a)} Let $t\in\{2,\ldots,m-1\}$ and
let for each $\bar\mu\in V^t$ the set
$B_{\bar\mu}$ be a ${\bar\mu}$-component of order $t-1$.
Then the set $B=\bigcup_{\bar\mu\in V^t} B_{\bar\mu}$ is a component of order $t$.

{\rm (b)} If $B$ is a component of order $t$ and $g^t\in {\cal A}^t$,
then $g^t(B)$ is a component of order $t$ too.

{\rm (c)} A component of order $m-1$ is an extended $1$-perfect code.
\end{proposition}
\begin{proof}
(a) follows from (\ref{eq:At}), the definition of a component of order $t$,
and the equality $\Omega(\bar\mu+A^{t-1})=\Omega(B_{\bar\mu})$,
which holds by the definition of $\bar\mu$-component of order $t-1$.

(b) and (c) are straightforward from the definitions.
\end{proof}

\section{ The LA Construction of Extended $1$-Perfect Codes}
Proposition \ref{pro:1} is all we need to see that the following construction
leads to an extended $1$-perfect code.
The idea of the construction is to take the Hamming code and apply
isometries of $F^n$ to parts of the code in such a way that the neighborhood of the part
does not change. We call such isometries \emph{local automorphisms}; a local automorphism
acts on a part of the code and does not change the neighborhood of that part.
At the first stage we take components of order $1$ as such parts;
at the second stage, components of order $2$; and so on.
At the last stage, we ``turn'' the whole code.

\begin{constr}[LA -- local automorphisms] \label{con:1}
Assume for any integer $t \in \{2,\ldots, m\}$ and for any words
${\bar r_{i}\in V^{i}}$, $i=t,\ldots,m-1$
we have a local automorphism $g_{\bar r_t,\ldots ,\bar r_{m-1}}\in {\cal A}^{t-1}$;
in particular, $g\in {\cal A}^{m-1}$.
Then (as follows by induction on $t$ from Proposition \ref{pro:1})
the set $C$ represented by the following formulas is an extended $1$-perfect code.
\begin{IEEEeqnarray}{rCl}
A^1_{\bar r_{2},\ldots ,\bar r_{m-1}} & \eqdf & V^1,\qquad \bar r_{i}\in V^i
\nonumber\\
A^{t}_{\bar r_{t+1},\ldots ,\bar r_{m-1}} & \eqdf & \bigcup_{\bar r_{t}\in V^{t}}
\left(\bar r_t+g_{\bar r_t,\ldots ,\bar r_{m-1}}(A^{t-1}_{\bar r_t,\ldots ,\bar r_{m-1}})\right),
\nonumber \\ \nonumber&&
 \qquad\qquad t=2,\ldots,m-1
\label{eq:2}
\\
C & \eqdf & g(A^{m-1}).
\end{IEEEeqnarray}
\end{constr}

In Construction \ref{con:1} each code can be obtained in more than one way.
To evaluate the number of the codes that can be constructed in this way we need
 stronger restrictions on local automorphisms.

Let
\begin{IEEEeqnarray*}{rCl}
{\cal B}^1 &\eqdf& {\rm Aut}(A^{1})
\\
{\cal B}^t &\eqdf& {\rm Aut}\big(\{\bar r+\Omega(A^{t-1})\}_{\bar r\in V^{t}}\big),\qquad t=2,\ldots ,m-1.
\end{IEEEeqnarray*}
For each $t=1,\ldots ,m-1$ we fix a set ${\cal D}^{t}$ of representatives of
cosets from ${\cal A}^{t}/{\cal B}^{t}$.
Moreover, we choose the representatives in such a way that the following holds:
for two cosets $D_1,D_2\in{\cal A}^{t}/{\cal B}^{t}$ and their representatives
$d_1,d_2\in {\cal D}^{t}$, $d_1 \in D_1$, $d_2\in D_2$, the equality
$D_1=\tau D_2$ with some translation $\tau$ implies $d_1=\tau d_2$
(this condition is essential for the definition of degenerate collection
and Proposition~\ref{p:bold} below).

It can be shown by induction that

\begin{proposition}
The restrictions
$
g_{\bar r_t,\ldots ,\bar r_{m-1}}\in {\cal D}^{t-1}
$
do not reduce the set of codes that can be represented by {\rm (\ref{eq:2})}.
\end{proposition}
\begin{proof}
Assume that
$$G^t \eqdf \bigcup_{\bar r \in V^t}\left(\bar r+g_{\bar r}(G^{t-1}_{\bar r})\right)$$
where $G^{t-1}_{\bar r}$ is a component of order $t-1$
and $g_{\bar r}\in{\cal A}^{t-1}$
for all $\bar r\in V^t$.
Let $g\in {\cal A}^t$ and $g=dh$ where $d\in{\cal D}^t$
and $h\in {\cal B}^t$.

We claim that
\begin{equation}\label{eq:cl}
    g (G^t) = d (G'^t) \quad \mbox{where} \quad G'^t \eqdf \bigcup_{\bar q \in V^t}\left(\bar q+g'_{\bar q}(G^{t-1}_{\rho\bar q})\right)
\end{equation}
for some $g'_{\bar q}\in{\cal A}^{t-1}$ and permutation $\rho:V^t\to V^t$.
Indeed, by the definition of ${\cal B}^t$, for all $\bar r\in V^t$ we have
$$
h(\bar r+\Omega(A^{t-1}))=\rho^{-1}\bar r+\Omega(A^{t-1})
$$
where $\rho$ is some permutation on $V^t$.
So, we see that $h_{\bar r}(\cdot)\eqdf \rho^{-1}\bar r+h(\bar r+\cdot)$
belongs to ${\cal A}^{t-1}$.
Then, replacing $\bar r$ by $\bar q\eqdf \rho^{-1}\bar r$,
we see that (\ref{eq:cl}) holds with
$g'_{\bar q} \eqdf h_{\rho\bar q} g_{\rho\bar q}$.

So, using (\ref{eq:cl}), we can step-by-step replace the operators
$g_{...}\in {\cal A}^t$ by  $d_{...}\in {\cal D}^t$,
starting from $t=m-1$ and finishing with $t=1$.
\end{proof}

Therefore the following construction gives the same set of codes
as Construction \ref{con:1}.
\begin{constr}[LA, upper bound] \label{con:2}
Assume that for any integer $t \in \{2,\ldots, m\}$ and for any words
${\bar r_{i}\in V^{i}}$, $i=t,\ldots,m-1$,
we have $g_{\bar r_t,\ldots ,\bar r_{m-1}}\in {\cal D}^{t-1}$;
in particular, $g\in {\cal D}^{m-1}$.
Then the set $C$ defined by the formulas {\rm (\ref{eq:2})} is an extended $1$-perfect code.
\end{constr}

As we will see below (Theorem~\ref{th:Kla}),
almost all ($n\to\infty$) codes represented by Construction \ref{con:2}
have a unique representation and this gives a good upper estimation
\begin{equation}\label{up}K_{\scriptscriptstyle LA}(n)\leq
|{\cal D}^{m-1}|\prod_{t=1}^{m-2}
|{\cal D}^t|^{|V_{t+1}|\cdot|V_{t+2}|\cdot\ldots \cdot|V_{m-1}|}
\end{equation}
for the number $K_{\scriptscriptstyle LA}(n)$ of different extended
$1$-perfect codes of length $n$
obtained by the method of local automorphisms (\emph{LA}), i.\,e., by
Construction \ref{con:1} or \ref{con:2}.
To show that the number of different LA codes is close to this value,
we need some more restrictions on $g_{\bar r_t,\ldots ,\bar r_{m-1}}$.

Assume $L$ is a linear subspace of $F^n$ and for each $\bar r\in L$ we have
$g_{\bar r}\in {\cal D}^{t}$ and
$g_{\bar r}(\cdot)= \bar v_{\bar r}+\pi_{\bar r}(\cdot)$.
We say that the collection $\{g_{\bar r}\}_{\bar r\in L}$ is \emph{degenerate}
{{if}} the following conditions hold:

$\bullet$ the permutation $\pi_{\bar r}$ does not depend on $\bar r$,
i.\,e., $\pi_{\bar r}=\pi$ for all $\bar r\in L$;

$\bullet$ the set $\{\bar r + \bar v_{\bar r} \,|\, {\bar r\in L} \}$ is an affine
subspace of $F^n$.

Otherwise we say that $\{g_{\bar r}\}_{\bar r\in L}$ is \emph{nondegenerate}.

\begin{constr}[LA, lower bound] \label{con:3}
In addition to the conditions of Construction \ref{con:2} we require that
the collection $\mathbf{g}_{\bar r_{t+1},\ldots ,\bar r_{m-1}}=\{g_{\bar r_t,\ldots ,\bar r_{m-1}}\in {\cal D}^{t-1}\}_{\bar r_t\in V^t}$
is nondegenerate for every
$t\in \{2,\ldots,{m-1}\}$, $\bar r_{t+1}\in  V^{t+1}$, \ldots, $\bar r_{m-1}\in  V^{m-1}$.
\end{constr}

\section{ Calculations}

In this section we establish some facts concerning
the structure of order-$t$ components and related objects,
on which the main result is based.
Given $G\subseteq \F$, put
$$\Theta(G) \eqdf \{\bar x\in \F\,|\, \Omega(\bar x)\subseteq \Omega(G)\};$$
clearly, $G\subseteq\Theta(G)$ and $\Omega(\Theta(G))=\Omega(G)$.
The following fact is also straightforward:
\begin{proposition} \label{p:Theta} For each $G,G'\subseteq \F$ the equality
$\Omega(G)=\Omega(G')$ means $\Theta(G)=\Theta(G')$ and vice~versa.
\end{proposition}
For each $t=1,\ldots,m$ and $\bar x=(\bar x_0,\ldots,\bar x_{2^t-1})\in (F^{2^{m-t}})^{2^t}=F^n$
define the generalized parity check
$$p^t(\bar x) \eqdf \sum_{i=0}^{2^t-1}\bar x_i.$$

\begin{proposition} \label{p:At}
Let $1\leq t \leq m-1$; then the following claims hold:\\
{\rm (a)} $p^t(\bar x)=\bar 0$ for all $\bar x\in A^t;$\\
{\rm (a')} $|A^t|=2^{2^{m-t}(2^t-1)-t};$\\
{\rm (b)} $\Omega(A^t)=\{\bar x\in F^n\,|\,wt(p^t(\bar x))=1\};$\\
{\rm (b')} $|\Omega(A^t)|=2^{2^{m-t}(2^t-1)+m-t};$\\
{\rm (c)} if $t<m-1$, then $\Theta(A^t)=\{\bar x\in F^n\,|\,p^t(\bar x)=\bar 0\};$\\
{\rm (c')} if $t<m-1$, then $|\Theta(A^t)|=2^{2^{m-t}(2^t-1)};$\\
{\rm (c'')} $\Theta(A^{m-1})=\F$.
\end{proposition}
\begin{proof}
(a) and (a') are straightforward from the definition of $A^t$.

(b') Since the code distance of $A^t$ is $4$, we have $|\Omega(A^t)|=n|A^t|$.

(b) It follows from (a) that $wt(p^t(\bar x))=1$ for all $\bar x\in\Omega(A^t)$.
On the other hand, by (b'), we have $|\Omega(A^t)|=|\{\bar x\in F^n\,|\,wt(p^t(\bar x))=1\}|$.

(c) It follows from (b) that $p^t(\bar x)=\bar 0$ implies $\bar x\in \Theta(A^t)$.
Assume $p^t(\bar x)\neq\bar 0$. If $t<m-1$, then there is
$\bar y\in \Omega(\bar x)$ such that $wt(p^t(\bar y))>1$; therefore, $\bar x\not\in\Theta(A^t)$.

(c') follows from (c).

(c'') It follows from (b) or (b') that $\Omega(A^{m-1})=\FF$; thus $\Theta(A^{m-1})=\F$.
\end{proof}

In what follows we will use the `array' representation of elements of $F^n$:
$$\bar x=(x^t_{0,0},...,x^t_{0,2^{m-t}-1},x^t_{1,0},
\ldots,
x^t_{2^{t}-1,2^{m-t}-1})=
(x^t_{i,j})_{i,j}$$
where indexes $i,j$ change in lexicographical order.
I.\,e., for each $t=1,\ldots,m-1$ an element $\bar x$ in $F^n$
can be viewed as $2^{t}\times 2^{m-t}$-array
$$
\left(
\begin{array}{l@{\ }l@{\ }l@{\ \ }l}
x^t_{0,0} & x^t_{0,1} & \ldots & x^t_{0,2^{m-t}-1}\\
\ldots &\ldots &\ldots &\ldots \\
x^t_{2^{t}-1,0}& x^t_{2^{t}-1,1}& \ldots & x^t_{2^{t}-1,2^{m-t}-1}
\end{array}
\right)
$$
In these terms, $p^t(\bar x)$ is the  sum of rows of $(x^t_{i,j})_{i,j}$.
For further calculations, we introduce the sets
\begin{IEEEeqnarray*}{rCl}
  {B}^1 & \eqdf & V^1  \\
  {B}^t & \eqdf & V^t + \Theta(A^{t-1}) 
  , \qquad t=2,\ldots,m-1
\end{IEEEeqnarray*}

\begin{proposition}\label{p:Bt} The sets ${B}^t$ satisfy the following properties: \\
{\rm (d)} $\displaystyle B^t=\{\bar x\in F^n\,|\,p^t(\bar x)=\bar 0 \mbox{ and } \sum_{j,\ {\rm even}\, i}x^t_{i,j}=0 \}$;
\hfill{\rm\refstepcounter{equation}(\theequation)\label{eq:Bt}}\\
{\rm (d')} $|B^t|=|\Theta(A^t)|/2$;\\
{\rm (d'')} ${\rm Aut}(B^t)={\cal B}^t$.
\end{proposition}
\begin{proof}
(d) and (d') are straightforward.
For $t=1$ the claim (d'') trivially holds. Assume $t>1$.
Using Proposition~\ref{p:Theta}, we get
${\cal B}^t ={\rm Aut}(\{\bar r+\Theta(A^{t-1})\}_{\bar r\in V^t})$.
Moreover,
\begin{IEEEeqnarray*}{rCl}
{\rm Aut}(\{\bar r+\Theta(A^{t-1})\}_{\bar r\in V^t})
&=& {\rm Aut}\bigg(\bigcup_{\bar r\in V^t}(\bar r+\Theta(A^{t-1}))\bigg)
 \\ & = & {\rm Aut}(B^{t})
\end{IEEEeqnarray*}
because, as follows from Proposition \ref{p:At}(c), the sets $\bar r+\Theta(A^{t-1})$ are
connected components of
distance-$2$ graph of $B^{t}$.
\end{proof}

\begin{proposition}\label{p:AtBt}
Let $1\leq t \leq m-1$; then the following claims hold:\\
{\rm (a)} if $t<m-1$, then
${\cal A}^t=({\cal P}^t\rightthreetimes {\cal Q}^t)\rightthreetimes{\cal R}^t$
where
\begin{itemize}
 \item for groups ${\cal G}$ and ${\cal G}'$, the notation
  ${\cal G}\rightthreetimes{\cal G}'$ means a semidirect product with
  a normal subgroup ${\cal G}'$;
    \item
 ${\cal P}^t\simeq S_{2^{m-t}}$ is the subgroup of column permutations
 $\psi:(x^t_{i,j})_{i,j}\to (x^t_{i,\psi(j)})_{i,j}$;
    \item
 ${\cal Q}^t\simeq (S_{2^{t}})^{2^{m-t}}$
 is the set of collections of permutations in every column
 $(\phi_0,\ldots,\phi_{2^{m-t}-1}):(x^t_{i,j})_{i,j} \to (x^t_{\phi_j(i),j})_{i,j}$;
    \item
 ${\cal R}^t\simeq Z_2^{2^{m-t}(2^{t}-1)}$
 is the set of translations
 $\bar z+$, $\bar z\in \Theta(A^t)$;
\end{itemize}
{\rm (a')} ${\cal A}^{m-1}\simeq S_n\rightthreetimes Z_2^{n-1}$;\\
{\rm (b)} if $t<m-1$, then
${\cal B}^t=({\cal P}^t\rightthreetimes \widehat{\cal Q}^t)\rightthreetimes\widehat{\cal R}^t$
where
\begin{itemize}
    \item $\widehat{\cal Q}^t\simeq (S_2\rightthreetimes(S_{2^{t-1}})^2)^{2^{m-t}}$,
     $\widehat{\cal Q}^t \subset {\cal Q}^t$;
    \item
 $\widehat{\cal R}^t\simeq Z_2^{2^{m-t}(2^{t}-1)-1}$
 is the set of translations $\tau_{\bar z}$, $\bar z\in V^t+\Theta(A^{t-1})$ where
 $\tau_{\bar z}(\bar x)\eqdf \bar z+\bar x$.
\end{itemize}
{\rm (b')} ${\cal B}^{m-1}={\cal A}^{m-2}\rightthreetimes \{\tau_{\bar 0},\tau_{(11110...0)}\}$.
\end{proposition}
\begin{proof}
(a)
First we observe that ${\cal A}^t ={\rm Aut}(\Theta(A^t))$.
Since, by Proposition \ref{p:At}(c), $\Theta(A^t)$ is linear,
it holds ${\cal A}^t= {\cal O}^t \rightthreetimes {\cal R}^t$ where
${\cal O}^t\subset {\cal A}^t$ consists of coordinate permutations
and ${\cal R}^t\subset {\cal A}^t$ is a group of translations.

 It follows from Proposition \ref{p:At}(c) that ${\cal O}^t$ consists
of the permutations that do not break columns, i.\,e.,
an admissible permutation permutes columns and permutes elements in each column.

(a') follows from Proposition \ref{p:At}(c'').

(b) By Proposition~\ref{p:Bt}(d''), we have
${\cal B}^t = {\rm Aut}(B^{t})$.
Since $B^t$ is linear,
it holds ${\cal B}^t= \widehat{\cal O}^t \rightthreetimes \widehat{\cal R}^t$
where
$\widehat{\cal O}^t\subset {\cal B}^t$ is the coordinate permutation subgroup
and $\widehat{\cal R}^t\subset {\cal B}^t$ is the translation subgroup of ${\cal B}^t$.

Using Proposition~\ref{p:Bt}(d), we see that an arbitrary permutation from $\widehat{\cal O}^t$
does not break the columns of $(x^t_{i,j})_{i,j}$ and, moreover,
in each column the permutation does not change the parity of row-indexes
or changes the parity of all row-indexes.
(Indeed, in the case $t<m-1$ for any other coordinate permutation $\pi$ we can find a weight $2$ or weight $4$
word $\bar x$ such that $\bar x$ satisfies (\ref{eq:Bt}) but $\pi\bar x$ does not.)
It can be directly checked that all such permutations belong to ${\cal B}^t$.

(b') In the case $t=m-1$ the group ${\cal B}^t$ contains some other permutations
and this case can be easily calculated directly.
\end{proof}

\begin{corollary}\label{p:Dt}
$|{\cal D}^{m-1}|=n!/6((n/4)!)^4$.
 If $t<m-1$, then
$$
|{\cal D}^t|
=2\left( {\frac {2^t!}{2(2^{t-1}!)^2}} \right)^{2^{m-t}}
=2\left( {\frac 12 \big\lgroup {{2^t}\atop{2^{t-1}}} \big\rgroup} \right)^{2^{m-t}}
.$$
In particular,
$|{\cal D}^1|=2$,
$|{\cal D}^2|=2\cdot 3^{\frac{n}{4}}$,
$|{\cal D}^3|=2\cdot 35^{\frac{n}{8}}$,
$|{\cal D}^4|=2\cdot 6435^{\frac{n}{16}}$.
\end{corollary}

We say that an order-$t$ component $G$ is \emph{bold} {{if}} ${<}G{>}=B^t$ where
${<}G{>}$ means the affine span of $G$ (i.\,e., the minimal affine subspace
including $G$; if $G\ni \bar 0$, then the affine span coincides with the linear span).

The next proposition
helps us to see that all codes given by Construction \ref{con:3} are
pairwise different.
\begin{proposition}\label{p:bold}
Let $1\leq t\leq m-1$;
for each $\bar r\in V^t$ let $G_{\bar r}$ be a bold order-$(t-1)$ component and
$g_{\bar r}\in {\cal D}^{t-1}$. Put
$$ G \eqdf \bigcup_{\bar r\in V^t}(\bar r+g_{\bar r}(G_{\bar r})).$$
Then \\
{\rm (a)} $G$ is bold if and only if the collection
$\{g_{\bar r}\}_{\bar r\in V^t}$ is nondegenerate;\\
{\rm (b)} if $g'_{\bar r}\in {\cal D}^{t-1}$ and $G'_{\bar r}$ is an order-$(t-1)$  component for all $\bar r\in V^t$ (it is not necessary to assume that $G'_{\bar r}$ are bold), then
$$ G = \bigcup_{\bar r\in V^t}(\bar r+g'_{\bar r}(G'_{\bar r}))$$
implies $G'_{\bar r}=G_{\bar r}$ and $g'_{\bar r}=g_{\bar r}$ for all $\bar r\in V^t$.
\end{proposition}
\begin{proof}
(a) By the definition of bold component we have ${<}G_{\bar r}{>}=B^{t-1}$; thus
\begin{IEEEeqnarray*}{rCl}
{<}G{>}&=&{<}\bigcup_{\bar r\in V^t}(\bar r+g_{\bar r}(G_{\bar r})){>}
\\
&=&
{<}\bigcup_{\bar r\in V^t}(\bar r+g_{\bar r}({<}G_{\bar r}{>})){>}
\\
&=&
{<}\bigcup_{\bar r\in V^t}(\bar r+g_{\bar r}(B^{t-1})){>}.
\end{IEEEeqnarray*}
Since $g_{\bar r}(B^{t-1})$ is a half of $\Theta(A^{t-1})$,
the affine span ${<}G{>}$ coincides either with
$\bigcup_{\bar r\in V^t}(\bar r+\Theta(A^{t-1}))=B^t$ (i.\,e., $G$ is bold)
or with $\bigcup_{\bar r\in V^t}(\bar r+g_{\bar r}(B^{t-1}))$ ($G$ is not bold).
It is clear that the last case occurs if and only if
the sets $g_{\bar r}(B^{t-1})$, $\bar r\in V^t$, are translations of each other
(i.\,e., $g_{\bar r}$ have a common coordinate permutation)
and the translation vectors compose an affine function on $V^t$.


(b) It suffices to show that
for arbitrary $g,g'\in{\cal D}^{t-1}$ and bold components $G_0$, $G'_0$ of order $t-1$
the inequality $g'\neq g$ implies $g'(G'_0)\neq g(G_0)$.
This holds because, by the definitions of ${\cal D}^{t-1}$ and bold components and
the fact that ${\cal B}^{t-1}={\rm Aut}(B^{t-1})$ (Proposition~\ref{p:Bt}(d'')),
$g'\neq g$ implies $g'({<}G'_0{>})\neq g({<}G_0{>})$.
\end{proof}

\begin{proposition}\label{p:DtVt} If $1\leq t<m-1$, then
the number of degenerate collections $\{g_{\bar r}\in {\cal D}^t\}_{\bar r\in V^{t+1}}$
is $|{\cal D}^t|\cdot|V^{t+1}|$.
\end{proposition}
\begin{proof}
Assume $1\leq t<m-1$. As follows from Proposition~\ref{p:AtBt} and the fact that
$\Theta(A^t)/B^t=2$ (Proposition~\ref{p:Bt}(d')),
for each coordinate permutation $\pi$ there are $2$ or $0$ elements $\bar v$ such that
the automorphism $\bar v + \pi(\cdot)$ belongs to ${\cal D}^t$.
Thus we have:

1) The number of different coordinate permutations in ${\cal D}^t$ is $|{\cal D}^t|/2$.

2) For each admissible coordinate permutation $\pi$
the number of collections
$\{{\bar v_{\bar r}}+\pi(\cdot)\}_{\bar r\in V^{t+1}}$ of
automorphisms from ${\cal D}^t$ such that the set
$\{\bar r + \bar v_{\bar r} \,|\, {\bar r\in V^{t+1}} \}$
is an affine subspace equals
the number of two-value functions $f:V^{t+1}\to\{\bar v_1,\bar v_2\}$
satisfying
$f(\bar r_1)+f(\bar r_2)+f(\bar r_3)=f(\bar r_1+\bar r_2+\bar r_3)$
for any $\bar r_1,\bar r_2,\bar r_3\in V^{t+1}$,
i.\,e.,
 the number $2|V^{t+1}|$ of affine
$\{0,1\}$-value functions on $V^{t+1}$.

 By the definition of degenerate collection, the proposition is proved.
\end{proof}
\section{ A Lower Bound on the Number of $1$-Perfect Codes}
Denote by $\widetilde K_{\scriptscriptstyle LA}(n)$ the number of different
extended $1$-perfect codes given by Construction \ref{con:3}.

\begin{theorem}\label{th:Kla}
The extended $1$-perfect codes from Construction \ref{con:3}
are pairwise different. The number
of such codes equals
 \newcommand{\KKNK}{{\textstyle
    \big\lgroup\hspace{-0.7ex}  {{k} \atop {k{/}2}}  \hspace{-0.7ex}\big\rgroup
    }^{\hspace{-0.8ex}\frac{n}{k}}} 
 \newcommand{\DmMinusOne}{\frac{n!}{6\left({\frac n4}\mbox{\rm\large !}\right)^{4}}}

$$\begin{array}{l}
\widetilde K_{\scriptscriptstyle LA}(n)
=
|{\cal D}^{m-1}|\hspace{-0.5ex}\displaystyle\prod_{t=1}^{m-2}\hspace{-0.8ex}\left(
|{\cal D}^t|^{|V_{t+1}|}\hspace{-0.3ex}-\hspace{-0.3ex}|{\cal
D}^t|\hspace{-0.3ex}\cdot\hspace{-0.3ex}{|V_{t+1}|}\right)^{|V_{t+2}|\cdot\ldots \cdot|V_{m-1}|}
\nonumber\\
\qquad=
\DmMinusOne
\displaystyle\prod_{k=2,4,8,\ldots,\frac n4}
\Bigg(
\left( 2\cdot 2^{\vphantom{i}^{-\frac{n}{k}}} \KKNK \right)^
{\EExp[-0.5ex]{2k}}
\\
\hspace{30ex}-
\KKNK
\cdot{2^{\vphantom{i}^{-\frac{n}{2k}}}}
\Bigg)^{\EExpLog[-0.5ex]{2k}}
\end{array}$$
In particular, $\widetilde K_{\scriptscriptstyle LA}(16)=15 692 092 416 000 000$,
$\widetilde K_{\scriptscriptstyle LA}(32)\approx 2^{2363.79}$.
The following is the asymptotic formula for $\widetilde K_{\scriptscriptstyle LA}(n)$:
\begin{IEEEeqnarray}{rCl}
\widetilde K_{\scriptscriptstyle LA}(n)&\sim&|{\cal D}^{m-1}|\prod_{t=1}^{m-2}
|{\cal D}^t|^{|V_{t+1}|\cdot|V_{t+2}|\cdot\ldots \cdot|V_{m-1}|}
\nonumber\\&=&
\DmMinusOne
\prod_{k=2,4,8,\ldots,\frac n4}
\left(
{2}\cdot{2^{\vphantom{i}^{-\frac{n}{k}}}}
\KKNK
\right)^{\EExpLog[-0.5ex]{k}}
\label{eq:4}\\
&=&
                        \ExpHam{2}
\cdot
(   3^{\Eexp{4}}  \cdot \ExpHam{4} )
\nonumber\\&&
\qquad\qquad\,\,\,{}\cdot
(  35^{\Eexp{8}}  \cdot \ExpHam{8} )
\nonumber\\&&
\qquad\quad\,{}\cdot
(6435^{\Eexp{16}} \cdot \ExpHam{16} )
\nonumber\\&&
\,\cdot
\dots\cdot
\left(\left(
{\textstyle\frac 12{\big\lgroup\hspace{-0.7ex}}{{n/4}\atop{n/8}}
 {\hspace{-0.7ex}\big\rgroup}
}
\right)^{2^3}\cdot 2^{2^1}\right)
\cdot
\DmMinusOne
\nonumber
\end{IEEEeqnarray}
\end{theorem}
\begin{proof}
The number of ways to define an extended $1$-perfect code
using formulas (\ref{eq:2})
with restrictions of Construction~\ref{con:3}
can be easily calculated by Corollary~\ref{p:Dt} and
Proposition~\ref{p:DtVt}.
Proposition~\ref{p:bold} guarantees that
different local automorphisms give different codes.
\end{proof}
Since there is a one-to-one correspondence (deleting the last symbol) between extended
$1$-perfect codes and {\it $1$-perfect codes}, we have the following:

\begin{theorem}[A lower bound]
The number $B(n-1)$ of
$1$-perfect binary codes of length $n-1=2^m-1$
satisfies
\begin{equation}\label{eq:lbLA}
B(n-1)\geq \widetilde K_{\scriptscriptstyle LA}(n)
\end{equation}
where the exact expression and the asymptotic form for
$\widetilde K_{\scriptscriptstyle LA}(n)$
are given in
Theorem \ref{th:Kla}.
\end{theorem}

As we can see, the previous lower bound
$\ExpHam{2}\cdot(   3^{\Eexp{4}}  \cdot \ExpHam{4} )$
\cite{Kro:quas_l_b} consists of two multipliers ($k=2,4$)
of (\ref{eq:4}).


\begin{conjecture}\label{cj:} The lower bound {\rm (\ref{eq:lbLA})} is asymptotically tight,
i.\,e., (\ref{eq:4}) is the asymptotic number of $1$-perfect binary codes of length $n-1=2^m-1$.
\end{conjecture}

This conjecture is supported by our knowledge about $1$-perfect codes of small ranks, i.\,e., of rank $+1$ and of rank $+2$.
The \emph{rank} of the code is the dimension of its affine span;
we say that a $1$-perfect code of length $n-1$ is of rank $+p$
{{if}} its rank is $r_H+p$
where $r_H$ is the dimension of the linear $1$-perfect code (Hamming code) of corresponding length.
(The notion `affine span' means the same as `linear span'
if the code contains $\bar 0$, but the affine span is invariant for the code translations.)
We know that the LA construction gives almost all codes of rank $+1$
and almost all codes of rank $+2$ (and, of course, some other codes).
Moreover, if the affine span is fixed, then the number of $1$-perfect codes of rank $+1$ equals asymptotically
$$ \ExpHam{2} $$
and of rank $+2$,
$$
\ExpHam{2}\cdot
(   3^{\Eexp{4}}\cdot
    \ExpHam{4} )
    $$
(if we do not fix the affine span of code, then these values must be multiplied by the indexes
$n!/2^{n/2}(\frac{n}{2}-1)(\frac{n}{2}-2)\ldots(\frac{n}{2}-\frac{n}{4})$ and
$n!/24^{n/4}(\frac{n}{4}-1)(\frac{n}{4}-2)\ldots(\frac{n}{4}-\frac{n}{8})$ respectively).
This knowledge comes from the representation of $1$-perfect
binary codes of rank $+1$ and $+2$ \cite{AvgHedSol:class}
and the asymptotic number $3^{n+1}2^{2^n +1}(1+o(1))$
of $n$-ary quasigroups of order $4$ \cite{KroPot:Lyap,PotKro:asymp}.

\begin{remark}
All the codes given by Construction \ref{con:3} have the rank deficiency ${\rm RD}=2$
(the maximum rank of extended $1$-perfect binary codes of length $n\geq 16$ equals $n-1$, see \cite{EV:94}; so, the rank deficiency is defined as ${\rm RD}(C)\eqdf (n-1)-{\rm rank}(C)$)
and
(as follows from the bound
$\dim({\rm kernel}(C))\geq 2^{{\rm RD}(C)}$
for binary $1$-perfect codes of rank at least $+2$,
see \cite[Corollary 2.6]{PheVil:RankKernel})
the dimension of kernel at least $4$,
where ${\rm kernel}(C)\eqdf \{\bar k\,|\,C+\bar k=C\}$.
The last fact means that the construction gives at least
$$ \frac{\widetilde K_{\scriptscriptstyle LA}(n)}{n!2^{n-5}}$$
nonequivalent extended $1$-perfect binary codes of length $n$ and
$$ \frac{\widetilde K_{\scriptscriptstyle LA}(n)}{(n-1)!2^{n-5}}$$
nonequivalent $1$-perfect binary codes of length $n-1$,
where $n!2^{n-1}$ is the number of isometries of $\F$
and $(n-1)!2^{n-1}$ is the number of isometries of $F^{n-1}$.
\end{remark}

Yes, Conjecture~\ref{cj:} implies that almost all (extended) $1$-perfect codes have the
rank $n-3$, which is not full ($n-1$ for extended $1$-perfect codes of length $n$),
and even not fore-full ($n-2$).
This lacks support from the length-$16$ codes,
see \cite{ZZ:2006:l16r14} (the most part of codes has rank $14=n-2$, but the number of full-rank extended $1$-perfect codes of length $16$ is small indeed, see. \cite{OstPot:16},\cite{OPP:16}),
but the LA construction has not ``gathered power'' when $n=16$ ($m=4$).
Indeed, for $m=4$, among the three multipliers of (\ref{eq:4}), the first one ($t=1$)
is almost the same as the second ($t=2$), and the multiplier $n!/6((n/4)!)^4$
is the largest, while asymptotically the first multiplier is the most powerful one.
On the other hand, the fact that almost all codes have not full rank
can be expected. For example, this holds for $4$-ary distance $2$ MDS codes
($n$-ary qua\-si\-groups of order $4$, see \cite{PotKro:asymp,KroPot:n-ary});
the rank (over $Z_2^2$) has three different values for these codes
(rank $n-1$ for linear codes of length $n$, rank $n-\frac{1}{2}=\log_4|Z_2^{2n-1}|$ for `semilinear' codes, and rank $n$),
but the class with the middle rank value is the most powerful.
It is also notable that the number of nonequivalent order $16$
Steiner quadruple systems of full rank $15$ is smaller than of rank $14$ \cite{KOP}.
\pagebreak


\label{sec:biblio}
\providecommand\href[2]{#2} \providecommand\url[1]{\href{#1}{#1}}

\end{document}